x/3-x/2=6 x/3-x/2=6 x/3-x/2=6 x/3-x/2=6 x/3-x/2=6 x/3-x/2=6 x/3-x/2=6 x/3-x/2=6

x/3-x/2=6
x/3-x/2=6
x/3-x/2=6
x/3-x/2=6
x/3-x/2=6
x/3-x/2=6
x/3-x/2=6
x/3-x/2=6

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

x/3 – x/2 =6

(2x – 3x)/6 = 6

-x/6 = 6

-x = 36

x = – 36

Thanks